### v1

@@ 1 Introduction
FRBs last from microseconds to milliseconds, and are characterized by very high dispersion measures indicating cosmological distances, which implies that they are 10 to 12 orders of magnitude brighter than radio sources within the Milky Way. Also, they are very common across the sky, with a FRB expected once every ten seconds (Keane & Petroff, 2015; Champion et al., 2016). While early single-dish telescopes like Parkes and Arecibo could detect only a few FRBs, the CHIME telescope (CHIME/FRB Collaboration et al., 2018) with its much larger field of view has detected hundreds (CHIME/FRB Collaboration et al., 2021). Observation of additional signals from FRB121102 led to the understanding that FRBs can be of a repeating nature (Spitler et al., 2014, 2016), showing that they are not always cataclysmic events. While follow-up observations with single-dish telescopes were difficult, CHIME has reported that around 10% of FRBs detected have a repeating feature (Andersen et al., 2020, 2023), and a subset of those have periodic activity windows. Repeating FRBs have wide-ranging properties, peaking at different frequencies, with the presence of multiple components (e.g., Chatterjee et al., 2017; Tendulkar et al., 2017). In many cases, precise localization can be carried out, and in these cases, the studies of the intergalactic medium are possible.
The rapid time variation in FRBs indicates small intrinsic sizes. This would imply that FRB signals are spatially coherent, which would lead to interference effects in cases of gravitational lensing. Several authors (Eichler, 2017; Katz et al., 2020; Kader et al., 2022; Leung et al., 2022; Kumar & Beniamini, 2023) have considered the possible lensing of FRBs by black holes or other compact masses. These works also note a significant complication, which is frequency dependent plasma lensing by the interstellar medium in the host galaxy and in the Milky Way.
@@ 2.1 Images and magnification
  • Third is the collective lensing by stars within a strongly-lensing galaxy or cluster, which splits the individual images formed (the so-called macro-images) into many micro-images. Here $\cal T$ is of microsecond scale and $\theta_{E}$ is of micro-arcsecond scale. The phenomenon is of interest for sources that are small enough, and has been extensively studied for lensed quasars (see e.g., Vernardos et al., 2023). The scale of $\theta_{E}$ is too small to resolve the images, but changes in brightness resulting from transverse motion is observable. It is potentially even more interesting for FRBs, because of the smaller source size.
@@ 2.2 Wave optics in the eikonal regime
While geometric optics is a good enough approximation for most astronomical sources that are large enough so radiation from different parts of the source reaches the observer incoherently, wave optics applies to small and far away sources. This applies to gravitational waves (e.g., Dai et al., 2020). The case of FRBs is actually simpler, because of the much higher frequencies, enabling an eikonal approximation.
@@
[…]
@@ 2.3 Plasma scattering
Figure 1: Illustration of turbulent ISM. There is no intrinsic scale, but in the simulations the field size is the same as the microlensing fields in left panels in Figs. 24 below.
@@ 3 A simple simulation
We start with the illustrative microlensing simulation taken from Vernardos et al. (2023). This has a field of 50 point masses. To this, we add the plasma scattering term (15) with adjustable $\nu_{0}$. The micro-images are found using a simple recursive grid search. A random coherent signal spectrum $S(\nu)$ is generated and this signal is lensed according to Eqs. (1214). The lensed signal is then auto-correlated. Using Eqs. (1214), the auto-correlation turns out to be
[…]
@@ 3.1 Lensing-dominated regime
The observable signal is completely dominated by the few bright micro-images. The faint micro-images, though much more numerous, have a negligible contribution to the light.
@@ 3.2 The effect of plasma scattering
Figure 4: Arrival time contours (left) and auto-correlation signals (right) for low-frequency signals, $\nu/\nu_{0}$ being 7 (top row), 6 (middle row), 5 (bottom row) .
@@ 4 Conclusions
Based on the above simulation, we make the following interpretation. If a lensed FRB fits into a Fresnel zone and is not smeared too much by the contents of the interstellar or intergalactic medium — making it a coherent source — then a single macro-image could be used for recognizing the lens. The well-known fine time structure in FRBs indicates that they are sufficiently small to be coherent. Since most strongly lensing galaxies are elliptical, lensed FRBs would be exposed to the low end of plasma scattering.

### v2

@@ 1 Introduction
FRBs last from microseconds to milliseconds, and are characterized by very high dispersion measures indicating cosmological distances, which implies that they are 10 to 12 orders of magnitude brighter than radio sources within the Milky Way. Also, they are very common across the sky, with a FRB expected once every ten seconds (Keane and Petroff, 2015; Champion et al., 2016). While early single-dish telescopes like Parkes and Arecibo could detect only a few FRBs, the CHIME telescope (CHIME/FRB Collaboration et al., 2018) with its much larger field of view has detected hundreds (CHIME/FRB Collaboration et al., 2021). Observation of additional signals from FRB121102 led to the understanding that FRBs can be of a repeating nature (Spitler et al., 2014, 2016), showing that they are not always cataclysmic events. While follow-up observations with single-dish telescopes were difficult, CHIME has reported that around 10% of FRBs detected have a repeating feature (Andersen et al., 2020, 2023), and a subset of those have periodic activity windows. Repeating FRBs have wide-ranging properties, peaking at different frequencies, with the presence of multiple components (e.g., Chatterjee et al., 2017; Tendulkar et al., 2017). In many cases, precise localization can be carried out, and in these cases, the studies of the intergalactic medium are possible.
The rapid time variation in FRBs indicates small intrinsic sizes. This would imply that FRB signals are spatially coherent, which would lead to interference effects in cases of gravitational lensing. Several authors (Eichler, 2017; Katz et al., 2020; Kader et al., 2022; Leung et al., 2022; Kumar and Beniamini, 2023) have considered the possible lensing of FRBs by black holes or other compact masses. These works also note a significant complication, which is frequency dependent plasma lensing by the interstellar medium in the host galaxy and in the Milky Way.
@@ 2.1 Images and magnification
  • Third is the collective lensing by stars within a strongly-lensing galaxy or cluster, which splits the individual images formed (the so-called macro-images) into many micro-images. Here $\cal T$ is of microsecond scale and $\theta_{E}$ is of micro-arcsecond scale. The phenomenon is of interest for sources that are small enough, and has been extensively studied for lensed quasars (see e.g., Vernardos et al., 2024). The scale of $\theta_{E}$ is too small to resolve the images, but changes in brightness resulting from transverse motion is observable. It is potentially even more interesting for FRBs, because of the smaller source size.
@@
The deflection of light rays by large-scale gravitational lenses like galaxies or galaxy-clusters is described by the macrolensing potential, denoted by $\psi_{B}(\theta)$ in Eq. (8). The macrolensing potential governs the overall distortion of light rays originating from background sources. The deflection angle at angular position $\theta$ in the sky is given by the gradient of the potential, $\nabla\psi_{B}(\theta)$. A Taylor expansion of the deflection angle yields, at leading order, a constant term corresponding to an effective shift in the apparent source position from $\beta$. The next-order term, involving the second derivative of the potential (or the first derivative of the deflection angle), captures the local stretching and shearing of the arrival-time surfaces. Microlensing simulations (going back to Paczynski, 1986) generally assume constant stretching and shearing, neglecting higher-order effects. For simpler simulations designed for illustrative purposes (e.g., Vernardos et al., 2024) the background potential may be neglected altogether. Its omission does not qualitatively affect our primary results.
@@ 2.2 Wave optics in the eikonal regime
While geometric optics is a good enough approximation for most astronomical sources that are large enough so radiation from different parts of the source reaches the observer incoherently, wave optics applies to small and far away sources. This applies to gravitational waves (e.g., Dai et al., 2020; Mishra et al., 2021). The case of FRBs is actually simpler, because of the much higher frequencies, enabling an eikonal approximation.
@@
$\alpha_{k}=|\>\mu_{k}\>|^{1/2}\times\cases{1}&\mathrm{for\;minima}\\ i&\mathrm{for\;saddle\;points}\\ -1&\mathrm{for\;maxima}{}$ (14)
@@ 2.3 Plasma scattering
The grid-size for the simulation is chosen as $1024\times 1024$, with each point corresponding to one Einstein radius. The physical distance corresponding to the Einstein radius (Eq. 3) is given by
${\cal D}\,\theta_{E}=\sqrt{c\,\cal T\,D}$ (19)
and is of order milli-parsecs or light-days. The small-scale cutoff $l_{0}$ is roughly $10^{-2}{\,\cal D\,}\theta_{E}$ and the large-scale cutoff $L_{0}$ corresponds to $10{\,\cal D\,}\theta_{E}$. The region of interest where microlensing is assumed to take place lies between these two scales, and the turbulence is expected to be scale-free.
Figure 1: Illustration of time delay due to turbulent ISM, normalized according to Eq. (16). The spatial scale is the same as the microlensing fields in left panels in Figs. 24 below.
@@ 3 A simple simulation
We start with illustrative microlensing simulation from the recent literature (Vernardos et al., 2024, —see especially their Figs. 4 and 6). The simulation has 50 equal point-mass lenses, whose Einstein radii cover $\sim 40\%$ of the image plane ($\kappa_{*}\sim 0.4$ in the usual notation). This assume that the stellar mass is a significant contributor to the total lensing mass, but the stellar mass need not be dominant. More elaborate simulations have many thousands of particles, and include a smooth tidal field (known as external shear), but the qualitative properties of micro-images are already present in this example. Of particular interest is a cluster of four bright micro-images, two minima and two saddle points.
To the scaled time delay (6) corresponding to the above lensing system, we add the plasma scattering term (15) with adjustable $\nu_{0}$. The micro-images are found using a simple recursive grid search. A random coherent signal spectrum $S(\nu)$ is generated and this signal is lensed according to Eqs. (1214). The lensed signal is then auto-correlated. Using Eqs. (1214), the auto-correlation turns out to be
$C(t)$ $=\int S_{L}(t^{\prime})\,S_{L}^{*}(t^{\prime}-t)\,dt^{\prime}$ (20)
@@ 3.1 Lensing-dominated regime
An important property of collective microlensing by stars (see e.g., Nityananda and Ostriker, 1984; Saha and Williams, 2011) is that the observable signal is completely dominated by a few bright micro-images, while the faint micro-images, though much more numerous, have a negligible contribution to the light. In this example, the two bright minima and two bright saddle points all have $|\mu_{k}|\sim 1$ with the remaining images being orders of magnitude fainter. Thus only the first four micro-images would be observationally relevant. For a theoretical understanding, however, we will continue to consider the first few faint micro-images as well.
@@ 3.2 The effect of plasma scattering
Figure 4: Arrival time contours (left) and auto-correlation signals (right) for low-frequency signals, $\nu/\nu_{0}$ being 7 (top row), 6 (middle row), 5 (bottom row). The slight wiggliness of the curves in the bottom left panel are due to small-scale fluctuations from plasma scattering.
4 Observational aspects
Details of the observations and data processing go beyond the scope of this theoretically motivated manuscript, but we need to understand the chance of an actual detection of microlensing signatures.
FRBs are basically found as peaks in time series of the detected power over long observations. In reality, the data are channelised to form dynamic (time-resolved) spectra, to be able to correct for dispersion from the interstellar medium, and to distinguish astrophysical bursts from artificial interference. For the microlensing analysis, the original voltage signal $S(t)$, or equivalently $S(\nu)$, is needed for the duration of the burst, to form the autocorrelation according to Eq. (20). Such data are now recorded on a regular basis by CHIME and ASKAP, but also by telescopes like Effelsberg, which are used for follow-up observations of repeating bursts.
Because the central peak at $C(0)$ corresponds to the integrated detected power, we know it to be well above the noise level. Microlensing-induced peaks at $\tau_{k}-\tau_{j}$ (for $k>j\geq 0$) will be proportional to $|\alpha_{j}\alpha_{k}|$ or the geometric mean of the intensities of the corresponding micro-images. This can be compared to the central peak, which is effectively proportional to $\sum_{j}|\alpha_{j}|^{2}$ (the sum of all intensities), after averaging over frequency.
For the image configuration in Fig. 2, the ratio is 0.29 for the highest microlensing peak compared to $C(0)$, which is detectable if the initial FRB detection has a sufficient signal-to-noise ratio. The actual detection analysis requires a thorough statistical analysis of false-detection rates, taking into account the number of trial lags.
The frequency (here $\nu_{0}$) at which plasma scattering and lensing become comparable is unknown. If $\nu_{0}$ is lower than the GHz scale of FRB observations, the wave-optical effect discussed here will be observable. Otherwise the effect will be washed out by plasma scattering. The overall delay due to plasma scattering will be comparable to the observed dispersion (which is of order a second). This is vastly larger than the microsecond scale delays due to individual stars, but tiny compared to the overall time delay of to all the lensing mass (time delays in galaxy lensing are months). What is relevant is whether the plasma-scattering term varies on the microsecond scale over the transverse scale of microlensing (which is micro-arcseconds or picoradians). Equivalently, the question is whether the small-scale cutoff of the turbulent plasma is smaller than the microlensing scale.
The scattering delays can depend on the galactic latitude, with low galactic latitudes showing stronger effects in general (Wucknitz et al., 2021). For latitudes above $20^{\circ}$, a scattering delay at 1.4GHz is always less than 1 microsecond. Additional constraints on small-scale plasma fluctuations come from secondary-spectrum measurements (Sprenger et al., 2021, Fig. 4). Apart from a special case at 1 millisecond, most scattering features have delays of the order of a few 100 microseconds at 330MHz. As typical delays scale as $\sim\nu^{-4.4}$, even the strongest 1 millisecond feature would be $\sim 1.7$ microseconds at 1.4GHz - a case of strong scattering. Characteristic scattering angles at about 330MHz are also found to be $\sim 30$mas, which scale with frequency as $\nu^{-2.2}$, implying angular broadening of order $\sim 1$mas at 1.4GHz (Sprenger et al., 2021, Fig. 9). So, it turns out microseconds and mas are appropriate order-of-magnitude estimates at 1.4GHz. In the lensing galaxy however, scattering can be substantially stronger as the delays scale with distance. In such environments, plasma phase fluctuations may approach or exceed microlensing scales and could potentially suppress wave-optics features.
@@ 5 Conclusions
Based on the above simulation, we make the following interpretation. If a lensed FRB is smaller than the Fresnel scale — which is $\left({\cal D}\,c/\nu\right)^{1/2}\sim 10\,$au for GHz observations — and moreover is not smeared too much by the contents of the interstellar or intergalactic medium — making it a coherent source — then a single macro-image could be used for recognizing the lens. The well-known fine time structure in FRBs indicates that they are sufficiently small to be coherent. Since most strongly lensing galaxies are elliptical, lensed FRBs would be exposed to the low end of plasma scattering.